OpenStudy (anonymous):

Arthur is treating his friends to a meal at Chez Rich. The meals added up to $136.45. If he tips the waiter 15%, what is the amount of the tip. $20.46 $20.47 $13.36 $13.37

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (usukidoll):

15/100 x 136.45 = ?

OpenStudy (usukidoll):

we may need to round up on the cents... but it's not the last two choices.

OpenStudy (anonymous):

i picked A

OpenStudy (usukidoll):

20.4675 our number after 6 is greater than 5... round up on the cents

OpenStudy (anonymous):

so im correct

OpenStudy (usukidoll):

depends on what the question needs. if we don't have to round up then it's just 20.46 if we do have to round up thanks to that 7 after the 6 being greater than 5, then it's 20.47

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

On your last math test you earned an 88%. Of the 26 questions on the test, how many did you get correct? Round to the nearest whole number question. 12 22 22.88 23

OpenStudy (usukidoll):

88/100 x 26 = ?

OpenStudy (anonymous):

ok two more

OpenStudy (usukidoll):

sure but did you figure out the previous question?

OpenStudy (anonymous):

yess 22.88

OpenStudy (usukidoll):

correct

OpenStudy (anonymous):

87.5% of 48 = 135.5 42 39.5 420

OpenStudy (usukidoll):

87.5/100 x 48 = ?

OpenStudy (anonymous):

i got 42 for that

OpenStudy (usukidoll):

yup. I got 42 too.